OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (5, 4) and (x, 7).

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

answer is 8

OpenStudy (anonymous):

8 isn't a possible answer

OpenStudy (anonymous):

hold on im working it out

OpenStudy (anonymous):

i think its 5. but this is throwing me off

OpenStudy (anonymous):

hahaha, well thank you for trying :) I'll get some second opinions. @ocoombs12 @perl @hartnn Could either of you possibly help me with this? I've already missed like two questions on this assignment and if I miss another then I'll fail the assignment and have to retake it but it's taken my like thirty minutes to work it out the first time, haha.

OpenStudy (anonymous):

sorry..

OpenStudy (anonymous):

Don't apologize, it's alright :)

OpenStudy (anonymous):

aw okay thanks (: this isnt my strong suit

OpenStudy (anonymous):

Hahaha, mine either xD

OpenStudy (anonymous):

hehe right beautiful people dont need math (:

OpenStudy (anonymous):

@Secret-Ninja

hartnn (hartnn):

still need help with this ?

OpenStudy (anonymous):

No but I do need help with a similar question :)

hartnn (hartnn):

\(5*4 =7x\)

hartnn (hartnn):

same type \(2\times 8 =3x\)

OpenStudy (anonymous):

Oh, wait. With that thing you just provided.. Does that make the answer to my second question 16?

hartnn (hartnn):

2nd question ? for the last one you posted 16 =3x x = 16/3 = 5 1/3

OpenStudy (anonymous):

yeah, for the one with (2,8) and (x,3)

OpenStudy (anonymous):

ohhh, so it's 5 1/3? O.o

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

Thank you :)